# Learning in Structured Stackelberg Games

### v1

@@ 3.1 Leader’s task in the realizable setting
Consider the setting where the context space is $\pazocal{Z}=\{1,2\}\times[0,1]$ and we have 4 follower types $\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$. Follower types $f^{(1)}$ and $f^{(2)}$ have fixed utility functions $u_{f^{(1)}},u_{f^{(2)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ such that for all $\mathbf{z}\in\pazocal{Z}$, we have that $\pi_{*}^{(1)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, where $\pi_{*}^{(i)}$, refers to the optimal policy against follower $f^{(i)}$. Let follower type $f^{(3)}$ have utility function $u_{f^{(3)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$ and follower type $f^{(4)}$ have utility function $u_{f^{(4)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$. Let
$h_{\gamma_{1},\gamma_{2}}(\mathbf{z})=\begin{cases}f^{(1)}\;\;&\text{ if }0\leq\mathbf{z}[2]<\gamma_{1}\\ f^{(2)}\;\;&\text{ if }\gamma_{2}<\mathbf{z}[2]\leq 1\\ f^{(3)}\;\;&\text{ if }(\mathbf{z}[1]=1\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=2\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ f^{(4)}\;\;&\text{ if }(\mathbf{z}[1]=2\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=1\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ \end{cases}$
[…]
Figure 1: Follower type as a function of the context $\mathbf{z}$ in the example of Theorem 3.5. Observe that while $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the optimal mixed strategy to play is determined by whether $\mathbf{z}[2]\leq\frac{1}{2}$.
Theorem 3.5 shows that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information. The example constructed in our proof highlights that the Littlestone dimension is blind to how the hypothesis class interacts with the utilities of the players. This observation motivates the definition of a new quantity, which we term the Stakelberg-Littlestone (SL) dimension. In order to define the SL dimension, we first need to introduce the notion of an SL-Tree, which is a natural extension of the Littlestone tree to our setting.
Consider a tree of depth $d$ whose internal nodes are instances in $\pazocal{Z}$ and where the node-to-child edges for a single node, correspond to different labels in a finite set $\mathscr{K}$. The tree $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
Definition 3.6 ($\pazocal{H}$-Shattered Stackelberg-Littlestone Tree).
For a Stackelberg game, $\pazocal{G}$, let a Stackelberg-Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node, the node-to-child edges correspond to different vectors in $\Delta(\pazocal{A})$. We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ with respect to $\pazocal{G}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\mathscr{H}$, such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{h(\mathbf{z}_{i})}(\mathbf{z}_{i},\mathbf{x}))$. In the case of ties, we assume that the leader follows a fixed and arbitrary tie-breaking rule.
Equipped with these definitions, we are now ready to introduce the Stackelberg-Littlestone dimension, which characterizes the complexity of online learning in our setting.
Definition 3.7 (Stackelberg-Littlsestone Dimension).
The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary SL tree of depth $d$ that is shattered by $\pazocal{H}$.
3.1.1 Lower bound
We begin by showing that for any online Stackelberg game $\mathscr{G}$, no algorithm can achieve regret smaller than $\Omega(\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H}))$.
Theorem 3.8.
There exists a Stackelberg game, $\pazocal{G}$, such that for any hypothesis class $\pazocal{H}$ any deterministic algorithm will suffer at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret. Furthermore, any randomized algorithm will suffer at least $\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret in expectation.
Proof.
Consider a Stackelberg game, in which for any fixed context, playing the optimal strategy against the true follower type yields utility $1$, whereas playing any other strategy yields utility 00. Concretely, let $\beta_{1},\beta_{2},\ldots,\beta_{K}$ be distinct real numbers and consider the following payoff matrix, where the leader is the row player and the column player is follower type $f^{(i)}$. The leader has 2 pure actions and the follower has 3 pure actions.
$F_{1}$ $F_{2}$ $F_{3}$
$L_{1}$ $0,\frac{1}{\beta_{i}}$ $1,1$ $0,0$
$L_{2}$ $0,0$ $1,1$ $0,\frac{1}{1-\beta_{i}}$
Notice that if the leader puts probability $p>\beta_{i}$ on action $L_{1}$, then the follower will play action $F_{1}$ and the leader gets payoff 00. Similarly, if the leader puts probability $p<\beta_{i}$ on action $L_{1}$ then the follower plays action $F_{3}$ and the leader gets payoff 00. However, if the leader puts probability $p=\beta_{i}$ on action $L_{1}$, then the follower is indifferent and since the follower breaks ties in favor of the leader, she plays action $F_{2}$, giving the leader a payoff of $1$.
(Claim 1.) An adversary can force the learner to make $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by using the following strategy: (1) select $\mathbf{z}_{t}$ as the next input from an SL-Tree of depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ that is shattered by $\pazocal{H}$ and (2) set the optimal mixed strategy, $\mathbf{x}_{t}^{*}$ to be the edge in the shattered tree not chosen by the leader, that is $\mathbf{x}_{t}^{*}\neq\widehat{\mathbf{x}}_{t}$, where $\widehat{\mathbf{x}}_{t}$ is the strategy played by the learner at time $t$. Since the tree is shattered by $\pazocal{H}$, there must exist some $h\in\pazocal{H}$ that is consistent with $(\mathbf{z}_{t},f_{t})_{t=1}^{\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})}$. Therefore, any deterministic learner will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
3.1.2 Upper bounds
Having established that the Stackelberg-Littlestone dimension characterizes the difficulty of learning in our setting, in this section we present an algorithm and show that it achieves optimal regret. Our guarantees in this setting are with respect to the leader’s number of mistakes (i.e., timesteps where the utility of the leader is not optimal). By the realizability assumption and the boundedness of the utility function, all mistake bounds imply the same regret bounds.
Warm-up: context-free utilities.
As a warm-up to our main results, consider a simpler setting, in which both the leader and the follower utility functions do not take the context as an input (i.e. they are context-free). This means that the context at each timestep is now only predictive of the follower’s type at that timestep. In Theorem 3.9, we show that we can reduce the problem of learning the optimal policy of the leader to online multiclass prediction. Additionally, the known optimal algorithm in online multiclass prediction (the Standard Optimal Algorithm (SOA)) meets the lower bound shown in Theorem 3.8.
Theorem 3.9.
Consider a Stackelberg game, $\pazocal{G}$, such that the utilities of both the leader and the follower are context-free. Then for any finite class, $\pazocal{H}$, of hypotheses $h:\pazocal{Z}\to\mathscr{K}$, there exists a reduction of this setting to multiclass prediction that runs in $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|,|\pazocal{H}|))$ time and achieves regret at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Proof.
The reduction is as follows. First, compute the optimal strategy against each type of follower offline. This can be achieved using linear programming and requires $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|))$ runtime [16]. Then create a (possibly many-to-one) look-up table mapping each follower type $f^{(i)}$ to an optimal strategy, $\mathbf{x}_{*}^{(i)}$, for the leader against that follower. Compute $\pazocal{H}^{\prime}=\{h^{\prime}(\mathbf{z}):=\sum_{i=1}^{K}\mathbf{x}_{*}^{(i)}\cdot\mathbbm{1}\{h(\mathbf{z})=f^{(i)}\}\}_{h\in\pazocal{H}}$, which takes at most $O(\text{poly}(|\pazocal{H}|))$ time. Finally, we run the multiclass Standard Optimal Algorithm (see [17]) on $\pazocal{H}^{\prime}$ considering every optimal strategy for the leader in the look-up table as one label. It remains to show that $\textnormal{Ldim}(\pazocal{H}^{\prime})=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$. Notice that $\pazocal{H}^{\prime}$ is the set of all hypotheses, mapping each context to an optimal strategy for the leader against every follower type that is consistent with $\pazocal{H}$ for that context. Consider any SL tree and any Littlestone tree such that the context at each corresponding node is the same among them. If the SL tree is shattered by $\pazocal{H}$ then for every root-to-leaf path there must exist a hypothesis in $\pazocal{H}$ that is consistent with the labels along that path. But by definition of $\pazocal{H}^{\prime}$ for that corresponding path in the Littlestone tree there must exist an $h^{\prime}\in\pazocal{H}^{\prime}$ that is consistent with the labelings of that path. With the same argument, the converse is also true. Therefore, the SL tree with maximal depth that is shattered by $\pazocal{H}$ is precisely the generalized Littlestone tree with maximal depth that is shattered by $\pazocal{H}^{\prime}$. ∎
We note that the algorithm described above coincides with Algorithm 1 (presented below), which we prove to be optimal in the general contextual Stackelberg game setting.
Context-dependent utilities.
We now let the utilities of both players to be functions of the context. We present our main algorithm (Algorithm 1), which we show achieves regret that matches the lower bound in Theorem 3.8 for any Stackelberg game. Our approach is inspired by the Standard Optimal Algorithm for online multiclass prediction, with the main difference that our algorithm uses a hypothesis class, $\pazocal{H}$, to optimize directly over leader strategies in the simplex.
Intuitively, our algorithm keeps track of all the “surviving” hypotheses in $\pazocal{H}$, that is, the hypotheses that are consistent with the data seen so far. At each timestep the algorithm observes the context and considers the predicted follower types of all surviving hypotheses. For every such follower type the algorithm computes the optimal leader strategy against them. Then it plays the strategy that maximizes the Stackelberg Littlestone dimension of the set of surviving hypotheses were that to be the optimal strategy for that timestep. If the algorithm makes a mistake, then the set of surviving hypotheses will have strictly smaller Stackelberg-Littlestone dimension than before. In other words, the algorithm learns from this mistake. Notice that the number of possible strategies in the simplex is infinite, yet our algorithm only considers a bounded number of them at each timestep, at most one for each follower type.
@@ 3.1.2 Upper bounds
1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}$, where $h:\pazocal{Z}\to\mathscr{K}$ for every $h\in\pazocal{H}$.
2:Set $V_{0}=\pazocal{H}$
@@
5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
8:     Play action $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:     Observe $f_{t}$
10:     Set $V_{t}=\{h\in V_{t-1}:h(\mathbf{z}_{t})=f_{t}\}$
11:end
@@
For any Stackelberg game, $\pazocal{G}$ and every hypothesis class, $\pazocal{H}$, the regret of Algorithm 1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
@@
Let $T_{e}\subseteq[T]$ be the set of time steps on which SSOA errs and let $T_{c}=[T]\setminus T_{e}$ be the set of time steps on which SSOA correctly predicts the optimal strategy against the true follower type. Recall from Remark 3.2 that we can make $\sum_{t\in T_{c}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$ arbitrarily small.
Now from the boundedness of the utility function, we also know that $\sum_{t\in T_{e}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\leq|T_{e}|$. It suffices to show that for every time step $t\in T_{e}$ that SSOA makes a mistake, we have that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})-1$. For the sake of contradiction, assume that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})=\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$. Given that SSOA chooses a strategy $\mathbf{x}$ that maximizes $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$ and that SSOA made a mistake at this timestep, there must have been at least two distinct strategies $\mathbf{x},\mathbf{x}^{\prime}$ such that $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x}^{\prime})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t})$. However, this would mean that we could construct a full binary SL tree that is shattered by $\pazocal{H}$ and that has depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})+1$, which leads to a contradiction.
3.1.3 Relationship with Littlestone dimension

### v2

@@ 3.3 Stackelberg-Littlestone Dimension
We prove by induction on the depth $d$ of shattered trees that across $d$ timesteps the learner can be forced to incur regret at least as much as the weight of the tree’s root node. Since the Stackelberg-Littlestone dimension is defined as the supremum root node weight over all finite-depth trees, proving this suffices to show the desired claim. In case the supremum in the definition of $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ is not achieved by any finite tree, there will always be an additive error of $\epsilon$ that will get arbitrarily small as $d\to\infty$.
(Base Case) We consider a shattered tree of depth 111. Then
$\rho_{\varnothing}$ $=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$
$=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}u(\mathbf{z}_{\varnothing},\pi_{h^{*}}(\mathbf{z}_{\varnothing}),b_{f^{(i)}}(\mathbf{z}_{\varnothing},\pi_{h^{*}}(\mathbf{z}_{\varnothing})))-u(\mathbf{z}_{\varnothing},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{\varnothing},\mathbf{x}))\biggr{]}.$
So $\rho_{\varnothing}$ corresponds to the min-max value of the one-shot Stackelberg game and therefore, to a lower bound on the instantaneous regret of the learner.
(Inductive step) We assume that for any tree of depth $d$ that is shattered by $\pazocal{H}$ under $\pazocal{G}$, the weight of the root node lower bounds the learner’s cumulative regret across $d$ timesteps. We now consider a tree $\mathscr{T}_{d+1}$ of depth $d+1$ that is shattered by $\pazocal{H}$ under Stackelberg game $\pazocal{G}$. We will prove that for such a tree, the adversary can guarantee regret at least $\rho_{\varnothing}$, where $\mathbf{z}_{\varnothing}$ is the root node of this tree. By Definition˜3.6, we know that $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i}\biggr{]}$. Let $\mathbf{x}_{*}\in\Delta(\pazocal{A})$ and $i_{*}\in[K]$ be the respective arginf and argmax values in the above expression.
If the learner chooses to play $\mathbf{x}_{*}$ at this round, then by definition the adversary can guarantee instantaneous regret equal to $r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})$ by selecting the true follower type to be $i_{*}$ (which is realizable by $\pazocal{H}$ by definition of a shattered tree). If this happens, then by the inductive hypothesis the adversary can guarantee regret $\rho_{i_{*}}$ in the next $d$ timesteps and regret $\rho_{\varnothing}$ overall.
We will now prove the same even when the learner plays a strategy different from $\mathbf{x}_{*}$. Say that there exists $\tilde{\mathbf{x}}\in\Delta(\pazocal{A})$ such that the learner can guarantee instantaneous regret $r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})<r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})$, where $\tilde{i}=\textnormal{argmax}_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i)})+\rho_{i}\biggr{]}$. There are two cases.
(Case 1) Suppose $\tilde{i}=i_{*}$. Then we know that
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i}\biggr{]}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i_{*})})+\rho_{i_{*}}\biggr{]}.$
Since $\mathbf{x}_{*}$ achieves the infimum value
$r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i_{*})})+\rho_{i_{*}}\;\;\forall\tilde{\mathbf{x}}\in\Delta(\pazocal{A})$
$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i_{*})})\;\;\forall\tilde{\mathbf{x}}\in\Delta(\pazocal{A}),$
which leads to a contradiction.
(Case 2) Suppose $\tilde{i}\neq i_{*}$. Then we have that
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i})\biggr{]}\leq\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x}^{\prime},f^{(i)})+\rho_{i})\biggr{]}\;\;\forall\mathbf{x}^{\prime}\in\Delta(\pazocal{A})$
Taking $\mathbf{x}^{\prime}=\tilde{\mathbf{x}}$, we get
$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i})\biggr{]}$ $\leq\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i)})+\rho_{i})\biggr{]}$
$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})+\rho_{\tilde{i}}$
$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ $\leq\rho_{\tilde{i}}-\rho_{i_{*}}$
We denote $\ell:=r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ and by definition of $\tilde{\mathbf{x}}$ we know that $\ell>0$. We get that $\rho_{\tilde{i}}\geq\rho_{i_{*}}+\ell$. Now by the boundedness of the utility function, we know that if $\rho_{i_{*}}+\ell\geq d$, then the above inequality leads to a contradiction (the learner cannot incur strictly more than $d$ regret over $d$ timesteps).
Otherwise, we have a subtree with root node weight at least $\rho_{i_{*}}+\ell$ that is shattered by the induced hypothesis space $\pazocal{H}^{(\mathbf{z}_{\varnothing}\to f^{(\tilde{i})})}$. By the inductive hypothesis we know that the adversary can force at least $\rho_{i_{*}}+\ell$ regret over the remaining $d$ timesteps and therefore at least $r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})+\rho_{i_{*}}+\ell=r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}=\rho_{\varnothing}$ regret over $d+1$ timesteps, as desired. ∎
3.4 Stackelberg Standard Optimal Algorithm
In this section, we present an algorithm that achieves instance-optimal regret that is equal to the Stackelberg-Littlestone dimension. We call Algorithm˜1 the Stackelberg Standard Optimal Algorithm (SSOA), which maintains a similar approach to the Standard Optimal Algorithm (SOA) from online multiclass prediction. Specifically, it keeps track of the subset of hypotheses that are consistent with the history so far (˜8) and plays the mixed strategy that minimizes the worst-case possible instantaneous regret “regularized” by the Stackelberg-Littlestone dimension of the induced hypothesis class (˜6). The latter works as a proxy to the difficulty of the learning task of the leader that would remain after the current round is over.
@@ 3.4 Stackelberg Standard Optimal Algorithm
1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$
2:Set $V_{1}=\pazocal{H}$
@@
5:  Let $u_{*}^{(i,\mathbf{z}_{t})}:=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{t},\mathbf{x}))$ for all $i\in V_{t}(\mathbf{z}_{t})$
6:  Play $\mathbf{x}_{t}=\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in V_{t}(\mathbf{z}_{t})}\left[u_{*}^{(j,\mathbf{z}_{t})}-u(\mathbf{z}_{t},\mathbf{x},b_{f^{(j)}}(\mathbf{z}_{t},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}\left(V_{t}^{(\mathbf{z}_{t}\to j)}\right)\right]$
7:  Observe $f_{t}$
8:  Set $V_{t+1}=V_{t}^{(\mathbf{z}_{t}\to f_{t})}$
9:end
In Theorem˜3.9, we prove the optimality of our algorithm. Our proof follows a high-level idea similar to that of the Standard Optimal Algorithm (SOA) from online multiclass learning: with each mistake, the learner’s task becomes easier, as the Littlestone dimension decreases. Similarly here we show that when the learner incurs positive instantaneous regret, the Stackelberg Littlestone dimension decreases by at least this amount.
@@
For any Stackelberg game, $\pazocal{G}$, and every hypothesis class, $\pazocal{H}$, the regret of Algorithm˜1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
@@
We will show that at each timestep $t$, if the learner incurred instantaneous regret $r_{t}=r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$, then the the Stackelberg-Littlestone dimension must have decreased by at least $r_{t}$. In other words, $\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t})-r_{t}$, where $V_{t}$ is the set of hypotheses consistent with the history up to timestep $t$. We have that the regret at timestep $t$ is
$r_{t}$ $=u(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})})-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))+(\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to f_{t})})-\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to f_{t})}))$
$\leq\max_{i\in V_{t}(\mathbf{z}_{t})}\left[u(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})})-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))+\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to i)})\right]-\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to f_{t})})$
$=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in V_{t}(\mathbf{z}_{t})}\left[u(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})})-u(\mathbf{z}_{t},\mathbf{x},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to i)})\right]-\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$ (by definition of Algorithm 1)
$\leq\sup_{\mathbf{z}\in\pazocal{Z}}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in V_{t}(\mathbf{z})}\left[u(\mathbf{z},\mathbf{x}_{*}^{(i,\mathbf{z})},b_{f_{t}}(\mathbf{z},\mathbf{x}_{*}^{(i,\mathbf{z})})-u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}\to i)})\right]-\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$
$=\textnormal{SLdim}_{\pazocal{G}}(V_{t})-\textnormal{SLdim}_{\pazocal{G}}(V_{t+1}),$
where $\mathbf{x}_{*}^{(i,\mathbf{z})}$ is the leader’s optimal strategy against follower $f^{(i)}$ for context $\mathbf{z}$. ∎
3.5 Relationship with Littlestone dimension